OpenStudy (anonymous):

binge drinker male = 0.11 female = 0.12 not a binge drinker male = 0.32 female = 0.45 what is the probability that a randomly selected male student is not a binge drinker? Is this a complement problem? P (A) + P (not A) = 1? and so P (male is a binge drinker : 0.11) + P (male is not a binge drinker: 0.32) = 0.11+ 0.32 = .43 and 1 - .43 = .57 ?

jimthompson5910 (jim_thompson5910):

P(not a binge drinker| male) = P(not a binge drinker AND male)/P(male) P(not a binge drinker| male) = 0.32/(0.11+0.32) P(not a binge drinker| male) = 0.74418604651162

OpenStudy (anonymous):

so, it is not a complement problem...I am not familiar with your method...what rule are you using please? I don't get probability at all! thanks for your kind response

jimthompson5910 (jim_thompson5910):

here's one way to think of it

jimthompson5910 (jim_thompson5910):

let's say there are 100 people total

jimthompson5910 (jim_thompson5910):

11 of them would be male binge drinkers notice how 11/100 = 0.11, so that fits with the probability

jimthompson5910 (jim_thompson5910):

32 males would not be binge drinkers 32/100 = 0.32, so that fits too

jimthompson5910 (jim_thompson5910):

12 females would be binge drinkers 12/100 =0.12

jimthompson5910 (jim_thompson5910):

45 females would not be binge drinkers

jimthompson5910 (jim_thompson5910):

and 45/100 = 0.45 so it all fits

jimthompson5910 (jim_thompson5910):

how many males are there total (in this pool of 100 people)?

OpenStudy (anonymous):

total males in group = 43

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

sorry, how many males are NOT binge drinkers (of the 100)

OpenStudy (anonymous):

0.32 or 32/100 are not BD

jimthompson5910 (jim_thompson5910):

there are 32 males who do not binge drink so because there are 43 males total, the probability is 32/43 = 0.74418604651162

jimthompson5910 (jim_thompson5910):

we are only focusing on the males because it specifically states that you chose a male

OpenStudy (anonymous):

I did understand that part but was trying to find a rule that fit the question...I never really grasped the whole chapter on probabilities...and the class is moving too fast.. thanks for explaining...

jimthompson5910 (jim_thompson5910):

well it's using bayes theorem but I think it's better to think of it in this way

OpenStudy (anonymous):

never heard of bayes theorem...test tonight over the whole semester...hopefully he will have only 1-2 probability questions...and not the hard ones :) not that it might not be interesting to understand probabilities better...but my brain doesn't work like that....I envy people who really get math..

jimthompson5910 (jim_thompson5910):

this isn't a probability course?

OpenStudy (anonymous):

No, its an intro to the practice of statistics...a crunched down, short version just for nurses going on for Master's....this whole class has shaken my confidence ....I am an A student but not in this class...I think I could get it better if we had the whole semester instead of just a few weeks...

jimthompson5910 (jim_thompson5910):

i hate classes that are supposed to be 5 months but are jammed in like 3 weeks

jimthompson5910 (jim_thompson5910):

but I'm sure you'll make it

OpenStudy (anonymous):

yes, it didn't help that the instructor is short-tempered, easily frustrated when I ask him to clarify something...and there is no math lab/no tutors available...it has been the most miserable experience...but tonight is it...I'll pass, even if not with a better grade...I could take it again elsewhere and see what happens...uuhhh..Did I I really say that....Yikes! Thanks again for sharing your knowledge...

jimthompson5910 (jim_thompson5910):

I'm sure you'll be fine. You just need confidence. You're welcome